# Bayesian Deep Learning for Discrete Choice

### v1

@@ 8.3 Swiss Train Data
For the stated preference dataset on Swiss train route choice, we computed the VOTT for the two route alternatives, as well as the out-of-sample prediction accuracy based on the average of the SGLD latent utilities. While previous studies have estimated VOTT values, few have reported out-of-sample predictive accuracies. However, we found a recent implementation by [31] that compares the predictive performance of ICLV models with that of discrete choice models enhanced with graph neural networks (GNNs) and data augmentation strategies with generative AI.
In Figure 14, we present histograms of the average value of travel time savings (VOTT) across individuals in the sample, computed by averaging over SGLD samples, under various $\ell_{2}$ regularization settings. For all configurations, the number of layers and hidden units in both the embedding and nonlinear components was selected through cross-validation, following the same approach as in the NYC case study. Across all settings—including the case without $\ell_{2}$ regularization applied to the scale parameters of the nonlinearities (Figure 14(c))—our model produces average VOTT estimates that align with behavioral expectations, indicating a positive willingness to pay for a one-hour reduction in travel time. For reference, Vrtic and Axhausen [30] report VOTT values ranging from 11.9 CHF/h for leisure trips to 52.4 CHF/h for business trips, with an overall average of 17.1 CHF/h for the full sample, regardless of trip purpose.
@@
SOTA - 75.53888From the GNN-MNL composite model presented in [31]. The reported metric in this paper is most likely raw accuracy. In our study, we rely on weighted accuracy, making the performance difference even more significant.
@@ 9 Conclusion
To our knowledge, this paper is the first to explore Bayesian deep learning for discrete choice modeling—beyond the brief introduction we previously presented in [10]. By leveraging SGLD, we are able to provide credible intervals for economic quantities such as marginal rates of substitution (e.g., values of travel time), and to make predictions using distributions over parameter sets rather than point estimates, which are prone to overfitting when the model is underspecified. With our model design and two-step training procedure—which combines convergence to behaviorally sound hypotheses with posterior sampling over more complex representations—we effectively address key concerns raised by the discrete choice community regarding the application of deep learning: model architectures that lack behavioral interpretability, unstable parameter estimates, and the absence demonstrations on how to quantify and represent uncertainty.

### v2

@@ 8.3 Swiss Train Data
For the stated preference dataset on Swiss train route choice, we computed the VOTT for the two route alternatives, as well as the out-of-sample prediction accuracy based on the average of the SGLD latent utilities. While previous studies have estimated VOTT values, few have reported out-of-sample predictive accuracies. However, we found a recent implementation by [Li2024] that compares the predictive performance of ICLV models with that of discrete choice models enhanced with graph neural networks (GNNs) and data augmentation strategies with generative AI.
In Figure 16, we present histograms of the average value of travel time savings (VOTT) across individuals in the sample, computed by averaging over SGLD samples, under various $\ell_{2}$ regularization settings. For all configurations, the number of layers and hidden units in both the embedding and nonlinear components was selected through cross-validation, following the same approach as in the NYC case study. Across all settings—including the case without $\ell_{2}$ regularization applied to the scale parameters of the nonlinearities (Figure 16(c))—our model produces average VOTT estimates that align with behavioral expectations, indicating a positive willingness to pay for a one-hour reduction in travel time. For reference, vrtic2002impact report VOTT values ranging from 11.9 CHF/h for leisure trips to 52.4 CHF/h for business trips, with an overall average of 17.1 CHF/h for the full sample, regardless of trip purpose.
@@
SOTA - 75.53888From the GNN-MNL composite model presented in [Li2024]. The reported metric in this paper is most likely raw accuracy. In our study, we rely on weighted accuracy, making the performance difference even more significant.
@@ 9 Conclusion
To our knowledge, this paper is the first to explore Bayesian deep learning for discrete choice modeling—beyond the brief introduction we previously presented in [2025arXiv250309786V]. By leveraging SGLD, we are able to provide credible intervals for economic quantities such as marginal rates of substitution (e.g., values of travel time), and to make predictions using distributions over parameter sets rather than point estimates, which are prone to overfitting when the model is underspecified. With our model design and two-step training procedure—which combines convergence to behaviorally sound hypotheses with posterior sampling over more complex representations—we effectively address key concerns raised by the discrete choice community regarding the application of deep learning: model architectures that lack behavioral interpretability, unstable parameter estimates, and the absence demonstrations on how to quantify and represent uncertainty.